# From Language to Action: A Review of Large Language Models as Autonomous Agents and Tool Users

### v1

@@ 7 Reasoning, planning, and memory of LLM agents
This section reviews the core cognitive functions: reasoning, planning, and memory, within agents based on LLM. We compare how these capabilities are designed and utilized across single-agent and multi-agent systems, highlighting common strategies and architectural distinctions. An overview of the reasoning, planning, and memory techniques employed in LLM-based agents in single-agent, multiagent, and widely used categories is summarized in Table 8.
@@ 7.1.1 Application-specific reasoning techniques
Application-specific reasoning techniques are designed to handle personalized tasks more effectively by addressing the limitations of standard approaches in specialized contexts. They improve performance by optimizing for accuracy, interpretability, or scalability, where general methods may fall short. Such techniques enable agents to incorporate domain-specific knowledge, such as scientific logic, API tool usage, or multi-agent interaction dynamics, into their reasoning processes.
Single-agent reasoning techniques. A range of reasoning strategies has been designed for single agent settings where an agent acts, plans and self-assesses autonomously. TOOLLLM [52] employed a search-based decision tree based on depth first to navigate over 16,000 real-world APIs, highlighting the strength of search-based reasoning in tool-rich environments. EASYTOOL [53] advanced task decomposition techniques, breaking high-level goals into modular subproblems to improve step-wise execution. MATRIX [98] introduced simulation-based self-critique for introspective evaluation. ToolEmu [110] introduced risk-prompted reasoning, allowing GPT-4 to assess the risks of tool execution within a sandboxed setup. Formal frameworks such as Theorem-of-Thought [49] decompose reasoning into abductive, deductive, and inductive subagents coordinated through belief propagation and NLI-based edge scoring.
@@
Collectively, these single-agent innovations demonstrate how LLMs can reason, evaluate and adapt in isolation, often tackling complex tasks that demand high degrees of autonomy, safety, and explainability.
Multi-Agent Reasoning Techniques. Several frameworks demonstrate specialized reasoning innovations tailored to multi-agent settings, showcasing how LLM agents collectively coordinate, communicate and reason in complex environments.
For example, Prompt-Structured Strategic Reasoning [108] was introduced for multi-agent reinforcement learning, enabling agents to collaborate in cooperative games effectively. AGENTVERSE [35] supports collaborative reasoning of multiple agents through structured logical representations such as logic grid puzzles and Modular Grounded Symbolic Modules. Modeling belief states is essential in dynamic and partially observable domains; a Theory-of-Mind (ToM) inspired belief module [104] was developed to enhance coordination in multi-agent text-based games.
@@ 7.1.2 Widely-used reasoning techniques
Existing reasoning methods such as CoT, ReAct, Self-Reflection, and Self-Critique are crucial to enabling LLMs to perform complex cognitive tasks. These techniques support step-by-step problem solving, adaptive decision-making, and collaborative reasoning, contributing to their widespread adoption and effectiveness in improving LLM reasoning capabilities. CoT prompting is one of the most prominent approaches [40, 104, 45, 47, 25, 110, 49, 105, 74, 63, 59, 117, 133, 62, 73, 121, 79]. CoT facilitates step-by-step reasoning by decomposing complex problems into intermediate logical steps, demonstrating success in various domains, including code generation, healthcare, and robotics.
Another notable technique, ReAct [13, 113, 102, 112, 91] integrates reasoning with real-time task-specific actions, allowing agents to adapt dynamically to environmental feedback, a feature particularly beneficial in interactive and decision-intensive settings. Furthermore, the Self-Reflection and Self-Critique mechanisms [91, 19, 105, 60, 63, 117, 122, 13, 107, 133, 153, 111, 86, 83, 56] are robust and widely used for their ability to empower agents to assess their outputs, identify errors, and improve subsequent reasoning through iterative metacognitive processes. Self-Consistency methods [39, 54] improve reliability by generating multiple reasoning trajectories and selecting the most coherent response, with variants such as Trajectory Level Self-Consistency further improving the accuracy of planning.
@@ 7.2 Planning in LLM-based agents
7.2.1 Application-specific reasoning techniques
@@
Single-agent planning techniques. Single-agent frameworks emphasize self-sufficient decision making and autonomous planning that support independent agent functionality. Examples include TIMEARENA [91], which employs heuristic planning through temporal task decomposition to simulate realistic and time-sensitive environments. ReHAC [102] applies multistep planning grounded in Markov Decision Processes (MDP), allowing agents to navigate human task allocations using formal probabilistic frameworks. Cognitive agents using task stack-based planning [47] achieve structured goal decomposition, particularly useful in household or task-intensive domains.
@@
Multi-agent planning techniques. Several studies have proposed multiagent planning strategies to address the complexities of coordination, delegation, and collaboration between distributed agents. For example, ProAgent [25] enabled modular planning supported by validation mechanisms to facilitate robust collaboration. Then, in the recommendation system domain, a plan-first execution paradigm [87] was adopted to enhance the reliability of action through comprehensive preaction planning. DeCoAgent [37] utilized the decomposition of tasks based on prompts based on a few shots to enable decentralized agents to collaborate autonomously through smart contracts, offering a scalable framework for coordination.
Adaptive planning also appears in causal strategy learning systems like [126], which implement a two-phase structure, experimentation followed by exploitation, to allow agents to dynamically adjust strategies over time. AGENTBOARD [90] introduces a progress rate metric to monitor sub-goal completion across agents, thus supporting continuous evaluation in collaborative settings. Star Agents [44] adopt an evolutionary strategy planning to enhance data optimization through agent evolution. Structured workflows are demonstrated in UrbanKGent [86], where fixed pipeline planning ensures deterministic construction of knowledge graphs in a multi-agent environment. SMART-LLM [61] tackles multi-robot coordination through coalition formation planning, dynamically forming agent teams to achieve shared objectives. Human-in-the-loop workflow planning in healthcare simulations [83] enables structured scenario designs, combining multiple agents under expert-defined protocols for controlled execution.
@@ 7.2.2 Widely-used planning techniques
Planning methods are vital for agents to perform effectively across diverse tasks, with several robust techniques widely used due to their versatility across domains. Multistep planning enables agents to sequence actions over extended horizons, adapting dynamically based on intermediate outcomes and environmental feedback [102, 124, 61, 135]. Task decomposition planning breaks down complex goals into manageable subtasks, facilitating efficient resource allocation and parallel execution [120, 37, 53, 91]. ReAct-based planning integrates reasoning with action, allowing agents to adjust plans in real time according to environmental feedback [112, 91, 132]. Reflexion planning incorporates feedback loops and self-correction mechanisms, helping agents learn from errors and improve future performance [91, 132, 102]. Re-planning and dynamic adaptation techniques enable agents to revise strategies responsively when initial plans fail or conditions change unexpectedly [118, 122]. Furthermore, plan generation and evaluation approaches improve overall planning quality by creating and assessing multiple candidate plans prior to execution [97, 116]. Together, these techniques provide intelligent agents with the flexibility and resilience required to operate effectively in dynamic and uncertain environments, highlighting the centrality of adaptive planning in the advance of autonomous decision-making.
@@ 7.3.1 Application-specific memory techniques
Additional methods involve rolling logs for robots [103], previous tool outputs in context [117], negotiated ontologies to prevent repetition [64], and concatenated dialogue contexts for interactive planning [122]. These techniques enhance adaptability and long-term task performance in single-agent systems.
@@ 7.3.2 Widely-used memory techniques
Various memory techniques have been proposed to support LLM agents in different operational contexts, each addressing specific information retention and recall requirements. Context window memory is one of the most commonly used methods [68, 130, 62, 153, 86, 129, 132, 61, 118, 81, 70, 55]. Using the internal mechanisms of the transformer, it retains recent inputs and preserves local coherence throughout the prompt. This approach enables agents to maintain consistency during short interactions without relying on external systems. In contrast, conversational and dialogue history stores complete sequences of past interactions to support context-aware responses and maintain continuity over extended sessions [97, 65, 81, 120, 69, 83, 77].
More complex memory implementations include Reflexion-style memory systems, which integrate recent context with structured representations of prior outcomes, allowing agents to update behavior based on previous successes and failures [67, 102, 13, 111]. Furthermore, RAG methods connect agents to external knowledge sources during inference, enabling access to factual information [75, 37, 83, 55]. The components of the working memory and the scratchpad temporarily hold intermediate reasoning steps and enable sequential problem solving [67, 49, 124]. In parallel, long-term episodic memory models iteratively capture and organize agent experiences, supporting generalization of past situations and improving task adaptation [68, 123]. Specific systems apply short-term memory optimization techniques for immediate reactivity tasks to focus on rapid access to updated data while minimizing overhead [111, 81, 73, 78].
Several systems implement hybrid memory mechanisms, including models that merge local prompt-based memory with persistent indexed documents [70]. These designs reflect ongoing efforts to construct memory architectures that support immediate responsiveness and long-term behavioral consistency.
8 Impact of prompting, fine-tuning and memory augmentation
Although foundational LLMs have extensive general knowledge and strong reasoning skills, they are not automatically independent agents. Turning these general models into specialized, goal-driven agents that can see, plan and act well requires a variety of advanced techniques[37, 38, 102]. The performance, reliability and independence of LLM-based agents are highly dependent on three key areas of improvement: prompting, fine-tuning, and memory enhancement. These methods help developers shape the agent’s behavior, give it expertise in specific areas, and prepare it for long-term complex tasks [69, 78, 82].
@@ 8.1 Prompt engineering: a non-parametric approach to dynamic control and role delegation
Prompt engineering has become a key method for directing the behavior of LLM-based agents[83, 99, 101]. Unlike parametric approaches like fine-tuning, which change a model’s weights, prompt engineering works during inference time. It offers a simple way to define tasks, assign roles, and control agent behavior dynamically without needing a lot of computing power. This method takes advantage of the natural instruction-following and in-context learning abilities of LLMs to manage complex tasks[49]. This includes everything from individual agent reasoning to complex teamwork among multiple agents[39]. The flexibility of prompting makes it the main way to set goals, integrate tools, and simulate complex social interactions. At the single-agent level, prompt engineering plays a key role in defining an agent’s operational settings and reasoning style. For instance, Shinn et al.[13] propose Reflexion, a framework where agents reflect on task feedback and keep this text in memory to improve decision-making. This approach leads to significant performance improvements without changing model weights. This type of self-correction, guided by prompts, also works well in dynamic situations. Behari et al.[107] uses a Decision-Language Model (DLM) to understand natural language policy goals, suggest reward functions as code, and adjust them based on the results of simulations. However, the effectiveness of this control can be limited by context length and complexity. George et al.[101] notes that agents may struggle to use important information when faced with long histories and distracting "red herring" facts, even with CoT prompting. Moreover, prompting is the main way to enable agents to use external tools, which enhances their abilities beyond their existing knowledge. Yuan et al.[53] introduce the EASYTOOL framework, which changes extensive tool documentation into clear, simple instructions. This helps LLMs choose and perform the correct functions. This idea has been successfully applied in specialized fields. Bran et al.[56] illustrate this with ChemCrow, which combines 18 expert-designed chemistry tools that an LLM agent learns to use through well-designed prompts to independently plan and carry out complex syntheses. In the medical field, Goodell et al.[55] show that giving agents task-specific calculation tools like OpenMedCalc via prompting greatly reduces math errors. This concept extends to multimodal contexts, where Yang et al.[117] apply prompts in LLM-Grounder to break down language queries and manage visual grounding tools. Meanwhile, Song et al.[118] connect agent plans in LLM-Planner by updating prompts with lists of visually perceived objects. Prompt engineering ranges from controlling individual agents to managing entire "societies" of agents. Here, system-level prompts set roles, responsibilities, and communication protocols. Ni et al.[34] show this with the MechAgents framework, which creates teams of agents with roles such as planner, coder, and critic to solve mechanics problems. Chen et al.[35] propose AGENTVERSE to create expert roles and task instructions in a zero-shot manner. To manage long-context tasks, Zhang et al.[41] introduce the Chain-of-Agents (CoA) framework, where a manager agent combines contributions from multiple worker agents, each handling a part of the document. Likewise, Klein et al.[36] suggest the FLEET OF AGENTS (FOA) framework, which uses a particle filtering method where many agents explore a search space, guided by prompts to improve the cost-quality balance. This approach has proven very effective in simulating complex social and economic systems that resemble human behavior. Zhang et al.[19] build LLM agent societies with distinct traits and thinking patterns using only prompts. These societies display behaviors such as conformity and consensus. Zhao et al.[85] employ prompts in CompeteAI to create a virtual town filled with competing restaurant agents to examine market dynamics. Liu et al.[105] present the CogMir framework, which uses prompts to encode social variables and induce prosocial irrationality, showing that LLM agents can repeatedly mimic human cognitive biases. This multi-agent, prompt-driven method has strong applications in specific fields. In clinical triage, Lu et al.[38] use TRIAGEAGENT to simulate a multi-disciplinary team. Chen et al.[77] apply a similar method, MAC, for diagnosing rare diseases. In scientific discovery, Ghafarollahi et al.[69] organize agents in ProtAgents to design new proteins together. Barra et al.[83] creates a workflow to automate the production of healthcare simulation scenarios. While prompt engineering is a strong non-parametric tool, it also works well with parametric fine-tuning. Prompting often serves as a way to create high-quality, structured datasets for training smaller, specialized agents. For example, Qin et al.[52] use ChatGPT in their ToolLLM framework to generate a large dataset of instructions and API call solutions. This dataset is then used to fine-tune LLaMA into the effective ToolLLaMA. Similarly, Qian et al.[66] develop Mistral-Interact by fine-tuning Mistral-7B on a set of user dialogues created by prompting GPT-4. Pang et al.[98] use the MATRIX framework to generate alignment data for fine-tuning. This method, as Wu et al.[74] point out, allows the skills of large, proprietary models to be distilled into smaller, open-source agents. On the other hand, fine-tuning can help agents better follow complex prompts or perform specific skills. Yin et al.[45] demonstrate that fine-tuning agents with the LUMOS framework on unified, high-quality annotations improves their planning and grounding abilities. Wang et al.[97] fine-tune a 7B model in SOTOPIA-$\pi$ using data from prompted social interactions. This makes it match the social intelligence of a much larger GPT-4-based agent. This technique proves particularly useful in embodied AI. Yang et al.[123] show that an LLM expert, enhanced through reflective prompting, can generate distillation data to fine-tune a Vision-Language Model (VLM), teaching it to navigate and interact with a visual world. This shows that while prompting offers dynamic, real-time control, fine-tuning can embed more robust and specialized abilities into agents.

### v2

@@ 7 Reasoning, planning, and memory of LLM agents
This section reviews the core cognitive functions: reasoning, planning, and memory, within agents based on LLM. We compare how these capabilities are designed and utilized across single-agent and multi-agent systems, highlighting common strategies and architectural distinctions. An overview of the reasoning, planning, and memory techniques employed in LLM-based agents in single-agent, multi-agent, and widely used categories is summarized in Table S2 of the Supplementary Material.
@@ 7.1.1 Application-specific reasoning techniques
Application-specific reasoning techniques address the limitations of standard approaches in specialized contexts by incorporating domain-specific knowledge, such as scientific logic, API tool usage, or multi-agent interaction dynamics.
Single-agent reasoning techniques. A range of reasoning strategies has been designed for single-agent settings where an agent acts, plans, and self-assesses autonomously. TOOLLLM [50] employed depth-first search-based decision trees to navigate over 16,000 real-world APIs, highlighting the strength of search-based reasoning in tool-rich environments. This reasoning technique outperforms ReACT by approximately 81% in average pass rate. This result suggests that the primary bottleneck in LLM agent reasoning lies less in algorithmic complexity and more in the model’s capacity for coherent long-horizon planning. EASYTOOL [51] advanced task decomposition by breaking high-level goals into modular sub-problems to improve step-wise execution. It significantly enhances LLM performance by providing concise, structured tool instructions. Compared to methods like ReAct and CoT prompting, it achieves higher success rates with fewer errors, demonstrating superior reasoning and tool-utilization capabilities. MATRIX [94] introduced simulation-based self-critique for introspective evaluation. ToolEmu [106] introduced risk-prompted reasoning, implemented based on the ReAct framework, which allows GPT-4 to assess the risks of tool execution within a sandboxed setup. It allows systematic risk analysis of LLM agents and expands evaluation capability across various tools and testing configurations. Formal frameworks such as Theorem-of-Thought (ToTh) [47] decompose reasoning into abductive, deductive, and inductive subagents coordinated through belief propagation and NLI-based edge scoring. Compared to widely utilized techniques like CoT, CoT-Decoding, and Self-Consistency, ToTh achieved up to 29% higher accuracy on symbolic tasks and consistently strong performance across complex symbolic and numerical reasoning benchmarks.
@@
Collectively, these single-agent innovations demonstrate how LLMs can reason, evaluate, and adapt in isolation, often tackling complex tasks that demand high degrees of autonomy, safety, and explainability.
Multi-Agent Reasoning Techniques. Several frameworks demonstrate specialized reasoning innovations tailored to multi-agent settings, showcasing how LLM agents collectively coordinate, communicate, and reason in complex environments.
For example, Prompt-Structured Strategic Reasoning [104] was introduced for multi-agent RL, enabling agents to collaborate in cooperative games effectively. AGENTVERSE [34] supports collaborative reasoning of multiple agents through structured logical representations such as logic grid puzzles and Modular Grounded Symbolic Modules. It addresses issues such as erroneous feedback and incomplete task coverage. In experiments, GPT-4 agent groups consistently achieve stronger performance than single-agent and CoT approaches across various reasoning tasks. Modeling belief states is essential in dynamic and partially observable domains; a Theory-of-Mind (ToM) inspired belief module [100] was developed to enhance coordination in multi-agent text-based games.
@@ 7.1.2 Widely-used reasoning techniques
Existing reasoning methods, such as CoT, ReAct, Self-Reflection, and Self-Critique, are crucial to enabling LLMs to perform complex cognitive tasks. These techniques support step-by-step problem solving, adaptive decision-making, and collaborative reasoning, contributing to their widespread adoption and effectiveness in improving LLM reasoning capabilities. CoT prompting [39, 100, 43, 45, 25, 106, 47, 101, 71, 61, 57, 113, 127, 60, 70, 117, 76] is one of the most prominent approaches. It decomposes complex problems into intermediate logical steps, demonstrating success across the code generation, healthcare, and robotics domains.
Another notable technique, ReAct [13, 109, 98, 108, 87] integrates reasoning with real-time task-specific actions, allowing agents to adapt dynamically to environmental feedback. This feature is particularly beneficial in interactive and decision-intensive settings. Furthermore, the Self-Reflection and Self-Critique mechanisms [87, 19, 101, 58, 61, 113, 118, 13, 103, 127, 146, 107, 83, 80, 54] are robust and widely used for their ability to empower agents to assess their outputs, identify errors, and improve subsequent reasoning through iterative metacognitive processes. Self-consistency methods [38, 52] improve reliability by generating multiple reasoning trajectories and selecting the most coherent response. This process, with variants such as Trajectory Level Self-Consistency, further improves the accuracy of planning.
@@
Comparatively, CoT exhibits strong performance in structured, stepwise reasoning but is less effective for interactive or dynamic tasks. Conversely, ReAct augments adaptability through the integration of reasoning and action, although it is associated with increased implementation complexity. Self-reflection facilitates iterative optimization of outputs but remains contingent upon adequate contextual information. Accordingly, the selection of an appropriate reasoning methodology requires rigorous consideration of task attributes, interaction requirements, and computational constraints.
@@ 7.2 Planning in LLM-based agents
7.2.1 Application-specific planning techniques
@@
Single-agent planning techniques. Single-agent frameworks emphasize self-sufficient decision making and autonomous planning that support independent agent functionality. Examples include TIMEARENA [87], which employs heuristic planning through temporal task decomposition to simulate realistic and time-sensitive environments. By addressing the critical issue of coordinating overlapping tasks with limited resources, it enables the evaluation and improvement of language agents’ multitasking and temporal reasoning capabilities. ReHAC [98] applies multistep planning grounded in Markov Decision Processes (MDP), allowing agents to navigate human task allocations using formal probabilistic frameworks. Unlike traditional rule-based or prompt-driven planning models, it comprises RL to generalize across different tasks, datasets, and collaboration paradigms, reducing reliance on extensive human input or domain expertise. Cognitive agents using task stack planning [45] achieve structured goal decomposition, particularly useful in household or task-intensive domains.
@@
Multi-agent planning techniques. Several studies have proposed multi-agent planning strategies to address the complexities of coordination, delegation, and collaboration between distributed agents. For example, ProAgent [25] enabled modular planning supported by validation mechanisms to facilitate robust collaboration. This approach outperforms traditional learning-based agents, particularly in zero-shot coordination scenarios where such agents typically struggle. Then, in the recommendation system domain, a plan-first execution paradigm [84] was adopted to improve the reliability of action through comprehensive pre-action planning. DeCoAgent [36] utilized the decomposition of tasks based on prompts, based on a few shots, to enable decentralized agents to collaborate autonomously through smart contracts, offering a scalable framework for coordination.
Adaptive planning also appears in causal strategy learning systems like [122]. This implements a two-phase structure, experimentation followed by exploitation, to allow agents to dynamically adjust strategies over time. AGENTBOARD [23] introduces a progress rate metric to monitor sub-goal completion across agents, thus supporting continuous evaluation in collaborative settings. It gives a clearer insight into agent capabilities compared to conventional success-rate benchmarks. Star Agents [42] adopt an evolutionary strategy planning to enhance data optimization through agent evolution. Structured workflows are demonstrated in UrbanKGent [83], where fixed pipeline planning ensures deterministic construction of knowledge graphs in a multi-agent environment. SMART-LLM [59] tackles multi-robot coordination through coalition formation planning, dynamically forming agent teams to achieve shared objectives. Human-in-the-loop workflow planning in healthcare simulations [80] enables structured scenario designs, combining multiple agents under expert-defined protocols for controlled execution.
@@ 7.2.2 Widely-used planning techniques
Planning methods are vital for agents to perform effectively across diverse tasks, with several robust techniques widely used due to their versatility across domains. Multistep planning enables agents to sequence actions over extended horizons, adapting dynamically based on intermediate outcomes and environmental feedback [98, 120, 59, 129]. Task decomposition planning breaks down complex goals into manageable subtasks, facilitating efficient resource allocation and parallel execution [116, 36, 51, 87]. Its main advantage is the reduction of overall problem complexity, but performance can degrade when subtasks are highly interdependent, requiring careful sequential coordination to ensure correct execution. ReAct-based planning integrates reasoning with action, allowing agents to adjust plans in real time according to environmental feedback [108, 87, 64]. While it can improve adaptability, its effectiveness varies with task complexity and depends heavily on accurate environment modeling. Reflexion planning incorporates feedback loops and self-correction mechanisms, helping agents learn from errors and improve future performance [87, 64, 98]. Re-planning and dynamic adaptation techniques enable agents to revise strategies responsively when initial plans fail or conditions change unexpectedly [114, 118], which can improve robustness in uncertain environments, though frequent replanning may introduce computational overhead. Furthermore, plan generation and evaluation approaches improve overall planning quality by creating and assessing multiple candidate plans prior to execution [93, 112].
@@ 7.3.1 Application-specific memory techniques
Additional methods involve rolling logs for robots [99], previous tool output in context [113], negotiated ontologies to prevent repetition [62], and concatenated dialogue contexts for interactive planning [118]. These techniques improve adaptability and long-term task performance in single-agent systems.
@@ 7.3.2 Widely-used memory techniques
Various memory techniques have been proposed to support LLM agents in different operational contexts, each addressing specific information retention and recall requirements. Context window memory is one of the most commonly used methods [65, 125, 60, 146, 83, 124, 59, 114, 78, 67, 53]. Using the internal mechanisms of the transformer, it retains recent inputs and preserves local coherence throughout the prompt. This approach enables agents to maintain consistency during short interactions without relying on external systems. In contrast, conversational and dialogue history stores complete sequences of past interactions to support context-aware responses and maintain continuity over extended sessions [93, 22, 78, 116, 66, 80, 74].
More complex memory implementations include Reflexion-style memory systems, which integrate recent context with structured representations of prior outcomes, allowing agents to update behavior based on previous successes and failures [64, 98, 13, 107]. Furthermore, RAG methods connect agents to external knowledge sources during inference, enabling access to factual information [72, 36, 80, 53]. The components of the working memory and the scratchpad temporarily hold intermediate reasoning steps and enable sequential problem solving [47, 120]. In parallel, long-term episodic memory models iteratively capture and organize agent experiences, supporting generalization of past situations and improving task adaptation [65, 119]. Specific systems apply short-term memory optimization techniques for immediate reactivity tasks to focus on rapid access to updated data while minimizing overhead [107, 78, 70, 75]. Several systems implement hybrid memory mechanisms, including models that merge local prompt-based memory with persistent indexed documents [67]. These designs reflect ongoing efforts to construct memory architectures that support immediate responsiveness and long-term behavioral consistency.
Different memory mechanisms vary in effectiveness according to the task and the context of the agent. Context window memory suits short-term, reactive tasks, while conversational and episodic memory support extended interactions and continuity. Retrieval-augmented and hybrid systems improve access to external knowledge, enhancing reasoning in complex domains. In multi-agent settings, structured and dynamic memory enable coordinated planning. In general, memory designs are most effective when aligned with specific operational demands.
8 Impact of prompting, fine-tuning, and memory augmentation
Although foundational LLMs have extensive general knowledge and strong reasoning skills, they are not automatically independent agents. Turning these general models into specialized, goal-driven agents that can see, plan, and act well requires a variety of advanced techniques[36, 37, 98]. The performance, reliability, and independence of LLM-based agents are highly dependent on three key areas of improvement: prompting, fine-tuning, and memory enhancement. These methods help developers shape the agent’s behavior, give it expertise in specific areas, and prepare it for complex long-term tasks [66, 75, 79].
@@ 8.1 Prompt engineering: a non-parametric approach to dynamic control and role delegation
Prompt engineering has become a key method for directing the behavior of LLM-based agents [80, 95, 97]. Unlike parametric approaches like fine-tuning, which change a model’s weights, prompt engineering works during inference time. It offers a simple way to define tasks, assign roles, and control agent behavior dynamically without needing a lot of computing power. This method takes advantage of the natural instruction-following and in-context learning abilities of LLMs to manage complex tasks [47]. This includes everything from individual agent reasoning to complex teamwork among multiple agents [38]. The flexibility of prompting makes it the primary way to set goals, integrate tools, and simulate complex social interactions. At the single-agent level, prompt engineering plays a key role in defining an agent’s operational settings and reasoning style. For instance, Reflexion [13] demonstrates how prompts enable self-reflection without weight updates, achieving performance improvements through feedback-guided adjustments. This type of self-correction, guided by prompts, also works well in dynamic situations. Behari et al. [103] used a Decision-Language Model (DLM) to understand natural language policy goals, suggest reward functions as code, and adjust them based on the results of simulations. However, the effectiveness of this control can be limited by context length and complexity. George et al. [97] notes that agents may struggle to use important information when faced with long histories and distracting "red herring" facts, even with CoT prompting. Moreover, prompting is the main way to enable agents to use external tools, which enhances their abilities beyond their existing knowledge. EASYTOOL [51] converts extensive tool documentation into clear, simple instructions. This helps LLMs choose and perform the correct functions and has been successfully applied in specialized fields. Besides, ChemCrow [54] combines 18 expert-designed chemistry tools that an LLM agent learns to use through well-designed prompts to independently plan and carry out complex syntheses. In the medical field, Goodell et al. [53] show that giving agents task-specific calculation tools like OpenMedCalc via prompting greatly reduces math errors. These domain-specific designs, ChemCrow and OpenMedCalc, yield higher precision in their respective fields yet require substantial expert knowledge. Therefore, selecting a prompt-tool strategy depends on whether adaptability or domain performance is prioritized. This concept extends to multimodal contexts, where Yang et al. [113] apply prompts in LLM-Grounder to break down language queries and manage visual grounding tools. Meanwhile, Song et al. [114] connect agent plans in LLM-Planner by updating prompts with lists of visually perceived objects. Prompt engineering ranges from controlling individual agents to managing entire "societies" of agents. Here, system-level prompts set roles, responsibilities, and communication protocols. Ni et al. [33] show this with the MechAgents framework, which creates teams of agents with roles such as planner, coder, and critic to solve mechanics problems. Chen et al. [34] propose AGENTVERSE to create expert roles and task instructions in a zero-shot manner. To manage long-context tasks, Zhang et al. [40] introduce the Chain-of-Agents (CoA) framework, where a manager agent combines contributions from multiple worker agents, each handling a part of the document. Likewise, Klein et al. [35] suggest the FLEET OF AGENTS (FOA) framework, which uses a particle filtering method where many agents explore a search space, guided by prompts to improve the cost-quality balance. This approach has proven very effective in simulating complex social and economic systems that resemble human behavior. Zhang et al. [19] build LLM agent societies with distinct traits and thinking patterns using only prompts. These societies display behaviors such as conformity and consensus. Zhao et al. [82] employ prompts in CompeteAI to create a virtual town filled with competing restaurant agents to examine market dynamics. Liu et al. [101] present the CogMir framework, which uses prompts to encode social variables and induce prosocial irrationality, showing that LLM agents can repeatedly mimic human cognitive biases. This multi-agent, prompt-driven method has strong applications in specific fields. In clinical triage, Lu et al. [37] use TRIAGEAGENT to simulate a multi-disciplinary team. Chen et al. [74] apply a similar method, MAC, for diagnosing rare diseases. In scientific discovery, Ghafarollahi et al. [66] organize agents in ProtAgents to design new proteins together. Barra et al. [80] create a workflow to automate the production of healthcare simulation scenarios. While prompt engineering is a strong non-parametric tool, it also works well with parametric fine-tuning. Prompting often serves as a way to create high-quality, structured datasets for training smaller, specialized agents. For example, Qin et al. [50] use ChatGPT in their ToolLLM framework to generate a large dataset of instructions and API call solutions. This dataset is then used to fine-tune LLaMA into the effective ToolLLaMA. Similarly, Qian et al. [63] develop Mistral-Interact by fine-tuning Mistral-7B on a set of user dialogues created by prompting GPT-4. Pang et al. [94] use the MATRIX framework to generate alignment data for fine-tuning. This method, as Wu et al. [71] point out, allows the skills of large, proprietary models to be distilled into smaller, open-source agents. On the other hand, fine-tuning can help agents better follow complex prompts or perform specific skills. Yin et al. [43] demonstrate that fine-tuning agents with the LUMOS framework on unified, high-quality annotations improves their planning and grounding abilities. Wang et al. [93] fine-tune a 7B model in SOTOPIA-$\pi$ using data from prompted social interactions. This makes it match the social intelligence of a much larger GPT-4-based agent. This technique proves particularly useful in embodied AI. Yang et al. [119] show that an LLM expert, enhanced through reflective prompting, can generate distillation data to fine-tune a Vision-Language Model (VLM), teaching it to navigate and interact with a visual world. This shows that while prompting offers dynamic, real-time control, fine-tuning can embed more robust and specialized abilities into agents.